a gaseous mixture contains the three noble gases helium, argon, and krypton. the total pressure exerted by the mixture is 1.57 atm and the partial pressures of he and ar are 0.33 atm and 0.39 atm repectively what is the partial pressure of the kr in the mixture
1
Expert's answer
2020-10-09T13:47:39-0400

The total pressure exerted by the mixture is 1.57 atm


The partial pressures of He 0.33 atm


The partial pressure of Ar 0.39 atm


Therefore:


1.57atm = (0.33+ 0.39+ Kr)


1.57atm. = 0.72 + Kr


Kr. = ( 1.57 - 0.72)


Kr = 0.85 atm